OpenStudy (anonymous):

A coffee blend contains Sumatra beans that cost $5/lb and Kona beans that cost $13/lb. If the blend costs $10/lb, how much of each type of coffee is in 50 lb of the blend?

OpenStudy (anonymous):

lets introduce a variable, say \(x\) is the number of pounds of Kona, so since the total is 50 there must be \(50-x\) pounds of Sumatra then you are going to end up with 50 pounds of $10 blend for a total of \(50\times\$ 10=\$500\) the \(x\) pounds of Kona cost \(13x\) and the \(50-x\) pounds of Sumatra cost \(\$5(50-x)\) for a total of \(13x+5(50-x)\) since you know the total is 500 set \[13x+5(50-x)=500\] and solve for \(x\)

OpenStudy (anonymous):

if you are used to solving these as a system of equations you would write \[x+y=50\] \[13x+5y=500\] but you would then solve for \(y\) in the first equation and get \(y=50-x\) substitute in to the second equation and get \[13x+5(50-x)=500\] so there is really no difference in method